# Evaluating the Effect of Retrieval Augmentation on Social Biases

### v2

@@ Abstract
Tianhui Zhang1   Yi Zhou2  Danushka Bollegala1 1University of Liverpool 2Cardiff University Tianhui.Zhang@liverpool.ac.uk ZhouY131@cardiff.ac.ukdanushka@liverpool.ac.uk
@@
In the context of social biases, a stereotype is a widely held but oversimplified and fixed belief or assumption about the characteristics, attributes, or behaviours of members of a particular social group. Social groups can be further categorised into advantaged or disadvantaged. (Nangia et al., 2020) Advantaged groups refer to demographic groups that historically had greater access to resources, opportunities, power, or social privilege, whereas disadvantaged groups are those who have historically had discrimination, stereotypes, or unequal resource distributions. Anti-stereotypes are positive stereotypes held for the advantaged group.
We analyse the bias propagation using Bias Question Answering (BBQ) Parrish et al. (2022), a QA-structured benchmark that assesses social biases in LLMs, across gender, age, race and religion applying three retrieval methods. Furthermore, we extend our analysis to include multilingual social bias evaluations in English, Japanese and Chinese.
@@
    Across four types of social biases, retrieving documents labelled with stereotypes significantly amplifies bias in the final output, even for generator LLMs that initially exhibit low intrinsic biases.
@@
However, these foundational studies are primarily limited to English and often utilize template-based contexts or a single collection of texts, which may not fully capture the complexity of real-world documents. Our work extends these works in several important directions: (i) decoupling the effects of the three core RAG components (document collection, retriever, and generator); (ii) employing controllable corpora with explicit stereotypical and anti-stereotypical polarities to precisely measure any bias amplifications; and (iii) extending the evaluation to a multilingual context and question sets (e.g. English, Chinese, and Japanese) across a richer set of human-annotated social biases.
@@
The document collection $\mathcal{D}$ itself can contain stereotypical content. If there are many documents that express various levels of stereotypical social biases, then a subset of those documents can be retrieved even when the query does not explicitly mention any social biases and in return influence the generator to produce a biased response. We evaluate this by using document collections with controlled levels of stereotypical and anti-stereotypical content on four types of social biases (i.e. gender, age, race, religion) (§ 4.2) in the document collection using three benchmark datasets covering English, Japanese Yanaka et al. (2024) and Chinese Huang and Xiong (2024) languages in § 4.5.
@@
Following the QA-based social bias evaluation approach proposed by Parrish et al. (2022), we evaluate the social biases in a modular RAG system based on its ability to correctly answer questions without reflecting any unfair stereotypical biases.
Accuracy
measures the model’s performance. For ambiguous contexts ($\textrm{Acc}_{\textrm{a}}$), it is the fraction of “UNKNOWN” answers. For disambiguated contexts ($\textrm{Acc}_{\textrm{d}}$), it is the fraction of correct answers based on the provided evidence. However, accuracy does not indicate the directionality of the bias (i.e. stereotypically biased towards the advantaged group vs. anti-stereotypically biased towards the disadvantaged group).
Diff-Bias (Jin et al., 2024)
addresses this by measuring the difference between stereotypical and anti-stereotypical responses(see Appendix A for the definition). A score of zero indicates no bias, while positive or negative scores indicate bias towards advantaged or disadvantaged groups, respectively. We report Diff-Bias for both ambiguous and disambiguated contexts. Due to space constraints, full accuracy results are presented in Appendix F.
@@
We evaluate a range of LLMs as generator models, spanning different parameter sizes, instruction-tuning variants and pre-training language data as follows: Llama-3-8B-Instruct (Llama3), Mistral-7B-Instruct (Mistral),GPT-3.5-turbo (GPT-3.5), Qwen2.5-3B/7B/14B/72B (Qwen) base and instruction-tuned versions. We use OpenAI API for GPT-3.5-turbo and o4-mini, while the remainder of the models are downloaded from Hugging Face. [^1] https://huggingface.co For document retrieval, we consider three methods: (a) VectorIndex from LlamaIndex with 1536-dimensional OpenAI text-embedding-ada-002 embeddings, (b) BM25, a sparse retriever available in LlamaIndex, and (c) Contriever, a contrastively pre-trained dense retrieval system Izacard et al. (2021) that uses the facebook/contriever retrieval model.
@@
Compared to the w/o RAG baseline (gray bars), the stereo-set (red bars) amplifies bias towards the advantaged group across all four bias categories. Conversely, the anti-set (blue bars) reduces the Diff-Bias score, often pushing the biases in the opposite direction, which indicates that the bias is not merely mitigated but reversed towards the disadvantaged group. This shows the high sensitivity of biases in RAG to the external documents.
@@
Prior work has shown that social bias categories are correlated and often results in intersectional biases across categories Tan and Celis (2019); Lalor et al. (2022); Ma et al. (2023). This has important implications for RAG where the queries and documents cover different social bias categories, even when the targeted bias category has been filtered or is absent from the retrieved documents. We found that such intersectional biases are also consistently amplified during RAG (see Appendix D), which raises serious concerns.
@@
While our main findings hold true on average across models (Figure 3), the choice of generator LLM introduces significant difference. We investigate three key aspects: the impact of instruction tuning, model scale, and specialised architectures like reasoning models. For this focused analysis, we present results on gender bias from the English BBQ dataset (see Table 2).
@@
To investigate our findings’ generalizability, we evaluate on Chinese (CBBQ (Huang and Xiong, 2024)) and Japanese (JBBQ (Yanaka et al., 2024)) datasets. As these languages lack dedicated stereotypical and anti-stereotypical corpora, we created document collections by machine-translating our English corpus. We also do human evaluation on the quality of translation in Appendix B.  [^2] Race and Religion categories are unavailable in JBBQ.
@@
The bias amplification trend for the ambiguous contexts is consistent for Japanese as seen from the results on JBBQ. Moreover, the analysis presented in Appendix B indicates that machine translation sometimes fail to preserve certain nuances of the original stereotypes, and Japanese-specific issues such as zero-pronoun resolution Isozaki and Hirao (2003) (i.e. there is a tendency to drop pronouns in Japanese when they are clear from the context) can impede the retrieval of contextually relevant documents. These findings show the importance of creating high-quality multilingual bias corpora.
@@
Vanilla 10.02 / -3.67 24.01* / 0.50 15.43 / -2.08 10.17 / -10.12
Table 5: Comparison of self-RAG against the vanilla base model Qwen-7B-Inst.. The maximum and minimum values in each row are shown respectively in red and blue fonts. 95 % CIs that do not overlap with the corresponding w/o RAG setting are indicated by *.
@@
While advanced RAG frameworks like Self-RAG (Asai et al., 2024) and RankRAG (Yu et al., 2024) often outperform standard RAG systems, their integrated and complex structure presents challenges for social bias analysis. It remains unclear whether they can mitigate biases in RAG.
The original Self-RAG model uses Llama-2 as the base model, which frequently refuses to answer social bias-related questions. This is problematic when evaluating social biases in RAG. Therefore, following the original training procedure and training data, we trained Self-RAG version using Qwen-7B-Inst. We then compare this Self-RAG version against the vanilla Qwen-7B-Inst. model.
Table 5 shows that in the w/o RAG setting, Self-RAG demonstrates a lower Diff-Bias score compared against the vanilla model, suggesting that its self-correction and adaptive retrieval mechanisms are effective at reducing the social biases in the baseline generator. However, when retrieving documents from the stereo-set, Self-RAG’s Diff-Bias score increases to 17.61, confirming that Self-RAG is also affected by the bias amplification in RAG. Furthermore, unlike the vanilla model, Self-RAG shows high Diff-Bias scores not only on the full-set but also on the anti-set, where the vanilla model’s bias is significantly lower.
@@ 4.7.1 Prompting Strategies
In-context Learning (ICL): Oba et al. (2024) used ICL to suppress biases in LLMs. We provide an LLM with eight curated few-shot examples from the English BBQ dataset to guide it towards unbiased generation (Oba et al., 2024). The examples cover ambiguous and disambiguated scenarios and are prepended to the user’s query (see Figure 5).
@@
As shown in Table 6, in the w/o RAG setting, both ICL and especially DDP are effective at mitigating the models’ intrinsic biases. For instance, DDP reduces the Diff-Bias of Qwen-7B-Inst. on ambiguous questions from 10.02 to nearly zero (0.69). However, in the w/o RAG setting, the Diff-Bias scores for ICL and DDP show little improvement over the default RAG baseline. This reveals a limitation of prompt-level strategies – the strong context injected by retrieved stereotypical documents can overwhelm prompt-based instructions or few-shot examples. In contrast, the summariser consistently reduces bias across all models in the RAG settings. Unlike other methods, the summariser directly tackles the biases from the retrieved documents before they influence the generator.
@@ 4.7.2 DPO Training
DPO (Rafailov et al., 2023) is an alignment algorithm that refines a model on preference data, using a “chosen” (preferred) and “rejected” response for each input. We apply DPO to investigate if reshaping a model’s preferences offers a more robust debiasing solution for RAG. Specifically, we fine-tune Qwen2.5‑7B‑Inst. on the GenderAlign dataset (Zhang et al., 2024), which contains 8000 dialogues paired with preferred (less biased) and rejected (more biased) responses.
Table 7 shows that DPO training successfully reduces bias in all conditions. The DPO-trained model exhibits a lower intrinsic bias in the w/o RAG setting. More importantly, it consistently achieves lower Diff-Bias scores across all retrieval sets (stereo, full, and anti) for ambiguous questions. The improvement on the disambiguated questions is smaller, which to be expected as the model is correctly relying on the provided context. This shows that DPO can be considered as a potential debiasing strategy for RAG systems by realigning the model’s internal behaviour.
Model w/o RAG stereo-set full-set anti-set
Vanilla 10.02 / -3.67 24.01 / 0.50 15.43 / -2.08 10.17 / -10.12
DPO 6.25 / -3.47 21.68 / 0.69 12.39 / -3.37 7.64 / -10.62
@@
We conducted a comprehensive study on how RAG influences social biases LLM on three languages and discovered that RAG amplifies social biases in LLMs when stereotypical documents collection are used for retrieval. We evaluate several potential mitigating strategies for RAG and show the effectiveness of DPO. We urge practitioners to move beyond evaluating LLMs in isolation, and consider a wholistic evaluation within RAG.
@@
Second, our analysis targeted three languages (i.e. English, Japanese, and Chinese) and four social bias types (i.e. gender, race, age, and religion). We selected those languages because of the availability of social bias evaluation datasets and the availability of native speakers of those languages who could evaluate the translation quality of the document collections. Furthermore, we selected those four bias types due to the availability of documents in the public domain that can be used in our retrieval experiments. The numbers of documents covering each social bias type is not equal as can be seen from Table 1. For example, most datasets cover gender-related social bias types well, whereas only four datasets include age-related social biases. Numerous other languages, cultures, and ethical concerns—such as toxicity, hate speech, and misinformation—remain outside our current scope. Evaluating RAG systems for these additional dimensions is a critical step for achieving broader safety and fairness guarantees.
@@
The document collections used for RAG are derived from publicly available sources as explained in Table 1, where each dataset’s original authors have annotated the documents for bias types. Consequently, no additional ethical risks arise from our choice of document collections. Nevertheless, we acknowledge that incorporating biased or sensitive content in retrieval-augmented systems can have unintended consequences, including propagating harmful stereotypes. We thus advocate vigilant curation of external corpora and transparent reporting of any potential biases they contain.
@@
Appendix A Evaluation Metrics
Accuracy:
${\rm Acc}_{a}=\frac{n_{au}}{n_{a}}$ (1)
${\rm Acc}_{d}=\frac{n_{ss}+n_{cc}}{n_{s}+n_{c}}$ (2)
Diff-bias Score:
$\text{Diff-bias}_{a}=\frac{n_{as}-n_{ac}}{n_{a}}$ (3)
$|\text{Diff-bias}_{a}|\leq 1-Acc_{a},\quad(0\leq Acc_{a}\leq 1)$ (4)
$\text{Diff-bias}_{d}=Acc_{ds}-Acc_{dc}=\frac{n_{ss}}{n_{s}}-\frac{n_{cc}}{n_{c}}$ (5)
$|\text{Diff-bias}_{d}|\leq 1-|2Acc_{d}-1|,\quad(0\leq Acc_{d}\leq 1)$ (6)
$=\begin{cases}2Acc_{d},&0\leq Acc_{d}\leq 0.5\\ 2(1-Acc_{d}),&0.5<Acc_{d}\leq 1\end{cases}$ (7)
Appendix B Translated Documents Human Examination

### v3

@@ Abstract
Tianhui Zhang1   Yi Zhou2  Danushka Bollegala1,3 1University of Liverpool 2Cardiff University 3Amazon Tianhui.Zhang@liverpool.ac.uk ZhouY131@cardiff.ac.ukdanushka@liverpool.ac.uk
@@
In the context of social biases, a stereotype is a widely held but oversimplified and fixed belief or assumption about the characteristics, attributes, or behaviours of members of a particular social group. Social groups can be further categorised into advantaged or disadvantaged (Nangia et al., 2020). Advantaged groups refer to demographic groups that historically had greater access to resources, opportunities, power, or social privilege, whereas disadvantaged groups are those who have historically had discrimination, stereotypes, or unequal resource distributions. Anti-stereotypes are positive stereotypes held for the advantaged group.
We analyse the bias propagation using Bias Question Answering (BBQ) Parrish et al. (2022), a QA-structured benchmark that assesses social biases in LLMs, across gender, age, race and religion applying three retrieval methods. Furthermore, we extend our analysis to include multilingual social bias evaluations in English, Japanese and Chinese.  [^1] The dataset and code are available here.
@@
    Across four social biases types, retrieving documents labelled with stereotypes significantly amplifies bias in the final output of 16 models, even for generator LLMs that initially exhibit low intrinsic biases.
@@
However, these foundational studies are primarily limited to English and often utilize template-based contexts or a single collection of texts, which may not fully capture the complexity of real-world documents. Our work extends these works in several important directions: (i) decoupling the effects of the three core RAG components (document collection, retriever, and generator); (ii) using controllable corpora with explicit stereotypical and anti-stereotypical polarities to precisely measure any bias amplifications; and (iii) extending the evaluation to a multilingual context and question sets (e.g. English, Chinese, and Japanese) across a richer set of human-annotated social biases.
@@
The document collection $\mathcal{D}$ itself can contain stereotypical content. If there are many documents that express various levels of stereotypical social biases, then a subset of those documents can be retrieved even when the query does not explicitly mention any social biases and in return influence the generator to produce a biased response. We evaluate this by using document collections with controlled levels of stereotypical and anti-stereotypical content on four types of social biases (i.e. gender, age, race, religion) (§ 4.2) in the document collection using three benchmark datasets covering English (Parrish et al., 2022), Japanese Yanaka et al. (2024) and Chinese Huang and Xiong (2024) languages in § 4.5.
@@
Accuracy:
${\rm Acc}_{a}=\frac{n_{au}}{n_{a}}$ (1)
${\rm Acc}_{d}=\frac{n_{ss}+n_{cc}}{n_{s}+n_{c}}$ (2)
Diff-bias Score:
$\text{Diff-bias}_{a}=\frac{n_{as}-n_{ac}}{n_{a}}$ (3)
$|\text{Diff-bias}_{a}|\leq 1-Acc_{a},\quad(0\leq Acc_{a}\leq 1)$ (4)
$\text{Diff-bias}_{d}=Acc_{ds}-Acc_{dc}=\frac{n_{ss}}{n_{s}}-\frac{n_{cc}}{n_{c}}$ (5)
$|\text{Diff-bias}_{d}|\leq\begin{cases}2Acc_{d},&0\leq Acc_{d}\leq 0.5,\\ 2(1-Acc_{d}),&0.5<Acc_{d}\leq 1.\end{cases}$ (6)
@@
We evaluate a range of LLMs as generator models, spanning different parameter sizes, instruction-tuning variants and pre-training language data as follows: Llama-3-8B-Instruct (Llama3), Mistral-7B-Instruct (Mistral), GPT-3.5-turbo (GPT-3.5), Qwen2.5-3B/7B/14B/72B (Qwen) base and instruction-tuned versions. We use OpenAI API for GPT-3.5-turbo and o4-mini, while the remainder of the models are downloaded from Hugging face. For document retrieval, we consider three methods: (a) VectorIndex from LlamaIndex with 1536-dimensional OpenAI text-embedding-ada-002 embeddings, (b) BM25, a sparse retriever available in LlamaIndex, and (c) Contriever, a contrastively pre-trained dense retrieval system Izacard et al. (2021) that uses the facebook/contriever retrieval model.
@@
Compared to the w/o RAG baseline (gray bars), the stereo-set (red bars) amplifies bias towards the advantaged group across all four bias categories. We also provide qualitative examples demonstrating how retrieved stereotypical documents can directly flip model predictions in ambiguous scenario in Appendix F. Conversely, the anti-set (blue bars) reduces the Diff-Bias score compared with w/o RAG baseline, which indicates that the bias is not merely mitigated but reversed towards the disadvantaged group. Furthermore, the Diff-Bias scores consistently decrease when the proportion of stereotypical documents is reduced (see Appendix D). This shows the high sensitivity of biases in RAG to the external documents.
@@
Prior work has shown that social bias categories are correlated and often results in intersectional biases across categories Tan and Celis (2019); Lalor et al. (2022); Ma et al. (2023). This has important implications for RAG where the queries and documents cover different social bias categories, even when the targeted bias category has been filtered or is absent from the retrieved documents. We found that such intersectional biases are also consistently amplified during RAG (see Appendix C), which raises serious concerns.
@@
While our main findings hold true on average across models (Figure 3), the choice of generator LLM introduces significant differences. We investigate three key aspects: the impact of instruction tuning, model scale, and specialised architectures like reasoning models. For this focused analysis, we present results on gender bias from the English BBQ dataset (see Table 2).
@@
To investigate our findings’ generalisability, we evaluate on Chinese (CBBQ (Huang and Xiong, 2024)) and Japanese (JBBQ (Yanaka et al., 2024)) datasets. As these languages lack dedicated stereotypical and anti-stereotypical corpora, we created document collections by machine-translating our English corpus. We also do human evaluation on the quality of translation in Appendix A. [^3] Race and Religion categories are unavailable in JBBQ.
@@
The bias amplification trend for the ambiguous contexts is consistent for Japanese as seen from the results on JBBQ. Moreover, the analysis presented in Appendix A indicates that machine translation sometimes fail to preserve certain nuances of the original stereotypes, and Japanese-specific issues such as zero-pronoun resolution Isozaki and Hirao (2003) (i.e. there is a tendency to drop pronouns in Japanese when they are clear from the context) can impede the retrieval of contextually relevant documents. These findings show the importance of creating high-quality multilingual bias corpora.
@@
CRAG 10.02 / -3.67 30.01* / 2.68 18.21 / -0.5 12.06 / -5.16
Table 5: Comparison of self-RAG and CRAG against the vanilla model Qwen-7B-Inst.. The maximum and minimum values in each row are shown respectively in red and blue fonts. 95 % CIs that do not overlap with the corresponding w/o RAG setting are indicated by *.
@@
While advanced RAG frameworks like Self-RAG (Asai et al., 2024) and CRAG (Yan et al., 2024) often outperform standard RAG systems, their integrated and complex structure presents challenges for social bias analysis. It remains unclear whether they can mitigate biases in RAG.
We apply both CRAG and Self-RAG into the experiments on gender bias type. For CRAG, we use Llamaindex to implement it. When retrieving, if the retrieved documents are incorrect, CRAG will drop the current documents and retrieve from the collections again. To make a fair comparison, we do not allow the CRAG to use web search.
For Self-RAG, as the original model uses Llama-2 as the base model, which frequently refuses to answer social bias-related questions. This is problematic when evaluating social biases in RAG. Therefore, following the original training procedure and training data, we trained Self-RAG version using Qwen-7B-Inst. We then compare this Self-RAG version against the vanilla Qwen-7B-Inst. model.
Table 5 shows that in the w/o RAG setting, Self-RAG demonstrates a lower Diff-Bias score compared against the vanilla model, suggesting that its self-correction and adaptive retrieval mechanisms are effective at reducing the social biases in the baseline generator. However, when applying CRAG, the Diff-Bias score increases compared the vanilla model, showing that the CRAG is not effective in mitigating the social bias.
When retrieving documents from the stereo-set, both Self-RAG and CRAG ’s Diff-Bias score increase, confirming that even advanced RAGs is also affected by the bias amplification in RAG. Furthermore, unlike the vanilla model, both advanced RAG show high Diff-Bias scores not only on the full-set but also on the anti-set, where the vanilla model’s bias is significantly lower.
@@ 4.7.1 Prompting Strategies
In-context Learning (ICL): Oba et al. (2024) used ICL to suppress biases in LLMs. We provide an LLM with four curated few-shot examples (see Figure 12) from the English BBQ dataset to guide it towards unbiased generation. The examples cover ambiguous and disambiguated scenarios and are prepended to the user’s query (see Figure 5).
@@
As shown in Table 6, in the w/o RAG setting, both ICL and especially DDP are effective at mitigating the models’ intrinsic biases. For instance, DDP reduces the Diff-Bias of Qwen-7B-Inst. on ambiguous questions from 10.02 to nearly zero (0.69). However, in the w/o RAG setting, the Diff-Bias scores for ICL and DDP show little improvement over the default RAG baseline. This reveals a limitation of prompt-level strategies – the strong context injected by retrieved stereotypical documents can overwhelm prompt-based instructions or few-shot examples. In contrast, the summariser consistently reduces bias across all models in the RAG settings. Unlike other methods, the summariser directly tackles the biases from the retrieved documents before they influence the generator. We also found these prompting-based debiasing methods do not harm factual accuracy on the BBQ tasks and some method such as Summariser can even improve it (§ G.6).
@@ 4.7.2 DPO Training
DPO (Rafailov et al., 2023) is an alignment algorithm that refines a model on preference data, using a “chosen” (preferred) and “rejected” response for each input. We apply DPO to investigate if reshaping a model’s preferences offers a more robust debiasing solution for RAG. Specifically, we fine-tune three LLMs (Qwen2.5‑7B‑Inst., Llama3-8B-Inst. and Qwen-14B-Inst) on the GenderAlign dataset (Zhang et al., 2024), which contains 8000 dialogues paired with preferred (less biased) and rejected (more biased) responses.
Table 7 shows that DPO training successfully reduces bias in all settings. The DPO-trained model exhibits a lower intrinsic bias in the w/o RAG setting. More importantly, it consistently achieves lower Diff-Bias scores across all retrieval stereo sets for ambiguous questions. The improvement on the disambiguated questions is smaller, which to be expected as the model is correctly relying on the provided context. This shows that DPO can be considered as a potential debiasing strategy for RAG systems by realigning the model’s internal behaviour.
Model Type w/o RAG stereo-set
Qwen-7B-Inst. Vanilla 10.02 / -3.67 24.01 / 0.50
DPO 6.25 / -3.47 21.68 / 0.69
Llama3-8B-Inst. Vanilla 5.65 / 1.69 14.68 / -0.40
DPO 3.42 / 0.49 10.06 / 0.38
Qwen-14B-Inst. Vanilla -2.38 / -2.38 4.61 / 2.68
DPO -1.78 / -5.06 3.32 / -0.99
@@
We conducted a comprehensive study on how RAG influences social biases LLM on three languages and 16 models discovered that RAG amplifies social biases in LLMs when stereotypical documents collection are used for retrieval. We evaluate several potential mitigating strategies including prompting strategies and DPO training for RAG and show the effectiveness of DPO in debiasing the effect of stereotypical documents. We urge practitioners to move beyond evaluating LLMs in isolation, and consider a wholistic evaluation within RAG.
Acknowledgements
Danushka Bollegala holds concurrent appointments as a Professor at University of Liverpool and as an Amazon Scholar. This paper describes work performed at the University of Liverpool and is not associated with Amazon.
@@ Acknowledgements
Second, our analysis targeted three languages (i.e. English, Japanese, and Chinese) and four social bias types (i.e. gender, race, age, and religion). We selected those languages because of the availability of social bias evaluation datasets and the availability of native of those languages within our research team who could evaluate the translation quality of the document collections. Furthermore, we selected those four bias types due to the availability of documents in the public domain that can be used in our retrieval experiments. The numbers of documents covering each social bias type is not the same as can be seen from Table 1. For example, most datasets cover gender-related social bias types well, whereas only four datasets include age-related social biases. Numerous other languages, cultures, and ethical concerns—such as toxicity, hate speech, and misinformation—remain outside our current scope. Evaluating RAG systems for these additional dimensions is a critical step for achieving broader safety and fairness guarantees.
@@
The document collections used for RAG are derived from publicly available sources as explained in Table 1, where each dataset’s original authors have annotated the documents for bias types. To our best knowledge, no ethical issues have been reported regarding those datasets. Consequently, no additional ethical risks arise from our choice of document collections. Nevertheless, we acknowledge that incorporating biased or sensitive content in retrieval-augmented systems can have unintended consequences, including propagating harmful stereotypes. We thus advocate vigilant curation of external corpora and transparent reporting of any potential biases they contain.
@@
Appendix A Translated Documents Human Examination